OpenStudy (anonymous):

need help will fan 2x4 – 54x

OpenStudy (uri):

2x4 – 54x 2x(x^3-27) ? @charlesbennett

OpenStudy (uri):

refresh your page.

OpenStudy (solomonzelman):

you need to integrate it, derive it, factor it, solve for root, ... what exactly do you want to be done with this, there are many many things that can be done with it.

OpenStudy (anonymous):

factor completely and simplify if possible

OpenStudy (solomonzelman):

Okay, so first, as uri already showed there is a common factor of 2x. When you factor out of 2x, you get 2x(x^3-27) But then you can use another rule, a^3-b^3=(a-b)(a^2 +ab + b^2) (hint: 27 = 3^3)

OpenStudy (anonymous):

okay thanks

OpenStudy (solomonzelman):

you welcome.